# references.bib

### v1

@@ 4 Experiments
WorldStrat [WorldStrat] 0.723 0.732 0.747 0.738 0.733 0.736
MSRResNet [MSRResNet] 0.743 0.739 0.733 0.794 0.725 0.783
DBPN [DBPN] 0.763 0.740 0.728 0.783 0.726 0.750
Pix2Pix [Pix2pix] 0.621 0.652 0.652 0.645 0.647 0.643
ControlNet [ControlNet] 0.625 0.653 0.648 0.650 0.658 0.644
DiffusionSat [DiffusionSat] 0.623 0.647 0.637 0.649 0.652 0.639
@@ 4.1 Datasets
We take the paired LR-HR satellite image dataset from [WorldStrat]. Each area of interest contains a single SPOT 6/7 high-resolution image with five bands. We take the RGB band of the SPOT6/7 satellite images, which has a GSD of 1.5 m/pixel. The low-resolution images are taken from the Sentinel-2 satellites consisting of 13 bands. We only pick the RGB band from them. For each area of interest, we have 16 paired low-resolution images taken at different time. The resolution ranges from 10 m/pixel to 60m per pixel. We crop the high-resolution image into 192x192 patches, and the low-resolution image into 63x63 patches.
@@
Function Map of the World (fMoW) [fmow] consists of high-resolution (GSD 0.3m-1.5m) satellite images of a variety of categories such as airports, amusement parks, crop fields and so on. However, its temporal resolution is limited due to its high resolution. We use the metadata of timestamp for pairing low-resolution Sentinel-2 images. Using the dataset provided in Cong et al. (2022), we create a fMoWSentinel-fMoW-RGB dataset with paired Sentinel-2 (10m-60m GSD) and fMoW (0.3-1.5m GSD) images at each of the original fMoW-RGB locations. Then, we use the bounding box provided by the metadata to extract relevant areas, and then crop the high-resolution images to patches of 512x512. For each high-resolution fMoW image, we find the corresponding Sentinel-2 images of the same location. We only take the RGB band of Sentinel-2 images and then apply the same cropping method as that for fMoW-RGB images.
@@ 4.2 Performance Benchmark
For super-resolution tasks on fMoW and WorldStrat datasets, we report perceptual quality metrics LPIPS to measure the perceptual similarity of the reconstructed image and the ground truth image, and FID to measure how realistic the reconstruction looks. We also report distortion metrics such as PSNR and SSIM for pixel-level similarity. Note that in satellite images, LPIPS is a more relevant metrics here as it measures the perceptual similarity, as mentioned in [DiffusionSat].
@@
WorldStrat [WorldStrat] 0.481 139.3 0.736 426.7
MSRResNet [MSRResNet] 0.472 159.7 0.783 286.5
DBPN [DBPN] 0.475 122.6 0.750 278.2
Pix2Pix [Pix2pix] 0.427 93.90 0.643 196.3
ControlNet [ControlNet] 0.580 108.0 0.644 102.3
DiffusionSat [DiffusionSat] 0.561 92.97 0.638 102.9
@@
We evaluate our algorithms on WorldStrat and fMoW LR-HR datasets. For each AOI (Area of Interest) of the WorldStrat dataset, we resize each LR image to $192\times 192$. We fix the prompt to be "Satellite Images" and compute the time difference between the low-resolution image and high-resolution image. Then, we add the time difference embedding network to the stable diffusion 1.2 and then fine-tune the stable diffusion 1.2 on the low-resolution and high-resolution pair of the WorldStrat dataset conditioning on the time difference. For each high-resolution image, we randomly select a low-resolution image from its 16 corresponding low-resolution images for training. After that, we get a latent diffusion model that takes a low-resolution image and the relative time difference as input and output the predicted high-resolution image.
For each AOI of the fMoW dataset, we resize each LR image to $512\times 512$ to align with the size of high-resolution image. We let the prompt to be "Satellite image of selected class from a direct overhead view", where selected class is the category name. We also use the same time difference embedding network and then fine tune the stable diffusion 1.2.
During inference, for both datsets we use 50 NFEs, and then perform optimization every 5 steps. More implementation details can be found in the supplementary materials.
For WorldStrat, we evaluate our algorithm on first 1000 images in the validation set. For fMoW, we select the first 100 images in the validation set from 6 classes: Airport, Amusement Pak, Car Dealership, Crop Field, Educational Institution, Electric Substation for evaluation.
@@
We consider six state-of-the-art baselines, WorldStrat [WorldStrat], MSRResNet [MSRResNet], DBPN [DBPN], Pix2Pix [Pix2pix], ControlNet [ControlNet], and DiffusionSat [DiffusionSat]. Both ControlNet and DiffusionSat are diffusion-based methods that takes LR as an additional condition for fine-tuned stable diffusion to predict the HR image. Both MSRResNet and DBPN are CNN-based methods that directly map the low-resolution image to the high-resolution image, and Pix2Pix is a GAN-based method. Both WorldStrat and DiffusionSat are fusion-based methods that fuse multiple low-resolution input to predict the HR image, while others take single low-resolution input and predict the HR image.
@@
ControlNet [ControlNet] 1427 12500 1427 21400
DiffusionSat [DiffusionSat] 1428 12500 1428 20000
@@
Number of images for fusion 2 4 6 8 10 12 14 16
LPIPS 0.41124 0.3967 0.3920 0.3891 0.3877 0.3874 0.3867 0.3866
Inference Time (s) 16.00 17.11 18.82 20.64 22.45 24.25 25.99 27.81
Table 5: Ablation study: number of images for fusion
@@ 5 Conclusion
Acknowledgments and Disclosure of Funding
The authors acknowledge support from Michigan Institute for Computational Discvoery and Engineering (MICDE) Catalyst Grant, and Michigan Institute for Data Science (MIDAS) PODS Grant.
Appendix A Appendix / supplemental material
@@ A.1 Implementation Details
WorldStrat We obtain the RGB satellite images provided by [WorldStrat]. We follow the preprocessing steps given by [WorldStrat], that crop the HR images into patches of $192\times 192$ and LR images into corresponding pathes of the HR images of size $63\times 63$. For each HR image, there are 16 corresponding LR images of different time. We extract the RGB band of the HR images and the RGB band of every LR images. We then resize LR image into the size of $192\times 192$. We use the same training and validation split provided by [WorldStrat] and then form a training and validation set. We then extract the timestamp from the metadata and compute the $dt_{i}$ for each $\text{LR}_{i}$.
fMoW We obtain the high-resolution images from [fmow], and the paired Sentinel-2 images from [satmae]. We first identify the area of interest on the HR images as given by [fmow], and then crop out other areas. Then we crop the corresponding LR images following the pre-processing steps given by [satmae]. We crop the HR images into patches of $512\times 512$, and align LR images into patches based on each HR image patch. Then we resize each LR image into the size of $512\times 512$ in accordance with the HR image. We consider 6 categories: airport, amusement parks, car dealership, crop field, educational institution, electric substation for training and testing. For training, we filter out HR images that do not have a corresponding LR, and those do not have three channels. We take the same training and validation split from [fmow]. When training, we consider all images from [fmow], and when testing, we pick the first 100 images from each selected category of the validation set of [fmow]. We also extract $dt_{i}$ from the metadata of fmow provided by [fmow], and the metadata of Sentinel-2 data provided by [satmae].
@@
We take the pretrained checkpoint (SD1.2) provided by [rombach2022highresolution], and then fine tune on the processed [fmow] and [WorldStrat] datasets. We rescale every LR image and HR image to the scale of [0,1]. Then, we use a learning rate $1e^{-5}$, and a batch size of 4 for both datasets. We stop training when the FID of sampled images stops improving. For [WorldStrat] dataset, we only train 800 iterations (partly due to the small dataset size). For [fmow] dataset, we train 4400 iterations. We then take the model for downstream inference tasks. For WorldStrat dataset, we use the prompt "Satellite images" for training. For fMoW dataset, we use the prompt "Satellite image of selected class from a direct overhead view", where selected class is the category name.
@@ A.3 Implementation Details of Baselines
We follow the original codebase of [WorldStrat], where we train the HighResNet model on both [fmow] and [WorldStrat] datasets. We tune the hyperparameters of the loss function based on our validation set. We stop training when the validation performance converges. On [WorldStrat] we train 125000 iterations with a batch size of 32, and on [fmow] we take 100000 iterations with a batch size of 32.
@@
We follow the original codebase of [MSRResNet], where we train the MSRResNet model on both [fmow] and [WorldStrat] datasets. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 160000 iterations for both [WorldStrat] and [fmow] datasets with a batch size of 16.
@@
We follow the original codebase of [DBPN], where we train the MSRResNet model on both [fmow] and [WorldStrat] datasets. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 100 epochs for both [WorldStrat] and [fmow] datasets with a batch size of 16.
@@
We follow the original codebase of [Pix2pix], where we train the MSRResNet model on both [fmow] and [WorldStrat] datasets. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 30 epochs for [fmow] and 100 epochs for [WorldStrat] with a batch size of 16.
@@
We follow the original codebase of [ControlNet]. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 12500 iterations with a batch size of 16 for [WorldStrat], and 21500 iterations with a batch size of 16 for [fmow].
@@
We implemented the 3D ControlNet architecture as mentioned in [DiffusionSat]. Then, we take the RGB band and the SWIR, NIR band from LR image for training 3D ControlNet. We tune the hyperparameters of the loss function based on validation set performance. We train for 12500 iterations with a batch size of 16 for [WorldStrat], and 20000 iterations with a batch size of 16 for [fmow]. We observe that further training worsened FID scores on both datasets.
@@ A.4 More Results
WorldStrat [WorldStrat] 17.98 0.396 13.42 0.443
MSRResNet [MSRResNet] 19.81 0.512 13.01 0.290
DBPN [DBPN] 19.17 0.471 11.90 0.268
Pix2Pix [Pix2pix] 19.76 0.448 12.21 0.180
ControlNet [ControlNet] 11.89 0.113 10.82 0.117
DiffusionSat [DiffusionSat] 12.34 0.133 10.63 0.109

### v2

@@ 4 Experiments
WorldStrat [Cornebise et al., 2022] 0.723 0.732 0.747 0.738 0.733 0.736
MSRResNet [Wang et al., 2018b] 0.743 0.739 0.733 0.794 0.725 0.783
DBPN [Haris et al., 2018] 0.763 0.740 0.728 0.783 0.726 0.750
Pix2Pix [Isola et al., 2017] 0.621 0.652 0.652 0.645 0.647 0.643
ControlNet [Zhang et al., 2023] 0.625 0.653 0.648 0.650 0.658 0.644
DiffusionSat [Khanna et al., 2024] 0.623 0.647 0.637 0.649 0.652 0.639
@@
12.52 0.122 0.565 102.3
15.28 0.272 0.496 105.1
17.40 0.396 0.418 88.12
Number of images for fusion 2 4 6 8 10 12 14 16
LPIPS 0.41124 0.3967 0.3920 0.3891 0.3877 0.3874 0.3867 0.3866
Inference Time (s) 16.00 17.11 18.82 20.64 22.45 24.25 25.99 27.81
Table 3: Ablation study: number of images for fusion
@@ 4.1 Datasets
We take the paired LR-HR satellite image dataset from [Cornebise et al., 2022]. Each area of interest contains a single SPOT 6/7 high-resolution image with five bands. We take the RGB band of the SPOT6/7 satellite images, which has a GSD of 1.5 m/pixel. The low-resolution images are taken from the Sentinel-2 satellites consisting of 13 bands. We only pick the RGB band from them. For each area of interest, we have 16 paired low-resolution images taken at different time. The resolution ranges from 10 m/pixel to 60m per pixel. We crop the high-resolution image into 192x192 patches, and the low-resolution image into 63x63 patches.
@@
Function Map of the World (fMoW) [Christie et al., 2018] consists of high-resolution (GSD 0.3m-1.5m) satellite images of a variety of categories such as airports, amusement parks, crop fields and so on. However, its temporal resolution is limited due to its high resolution. We use the metadata of timestamp for pairing low-resolution Sentinel-2 images. Using the dataset provided in Cong et al. (2022), we create a fMoWSentinel-fMoW-RGB dataset with paired Sentinel-2 (10m-60m GSD) and fMoW (0.3-1.5m GSD) images at each of the original fMoW-RGB locations. Then, we use the bounding box provided by the metadata to extract relevant areas, and then crop the high-resolution images to patches of 512x512. For each high-resolution fMoW image, we find the corresponding Sentinel-2 images of the same location. We only take the RGB band of Sentinel-2 images and then apply the same cropping method as that for fMoW-RGB images.
@@ 4.2 Performance Benchmark
For super-resolution tasks on fMoW and WorldStrat datasets, we report perceptual quality metrics LPIPS to measure the perceptual similarity of the reconstructed image and the ground truth image, and FID to measure how realistic the reconstruction looks. We also report distortion metrics such as PSNR and SSIM for pixel-level similarity. Note that in satellite images, LPIPS is a more relevant metrics here as it measures the perceptual similarity, as mentioned in [Khanna et al., 2024].
@@
WorldStrat 0.481 139.3 0.736 426.7
MSRResNet 0.472 159.7 0.783 286.5
DBPN 0.475 122.6 0.750 278.2
Pix2Pix 0.427 93.90 0.643 196.3
ControlNet 0.580 108.0 0.644 102.3
DiffusionSat 0.561 92.97 0.638 102.9
@@
We evaluate our algorithms on WorldStrat and fMoW LR-HR datasets. For each AOI (Area of Interest) of the WorldStrat dataset, we resize each LR image to $192\times 192$. We fix the prompt to be "Satellite Images" and compute the time difference between the low-resolution image and high-resolution image. Then, we add the time difference embedding network to the stable diffusion 1.2 and then fine-tune the stable diffusion 1.2 on the low-resolution and high-resolution pair of the WorldStrat dataset conditioning on the time difference. For each high-resolution image, we randomly select a low-resolution image from its 16 corresponding low-resolution images for training. After that, we get a latent diffusion model that takes a low-resolution image and the relative time difference as input and output the predicted high-resolution image. For each AOI of the fMoW dataset, we resize each LR image to $512\times 512$ to align with the size of high-resolution image. We let the prompt to be "Satellite image of selected class from a direct overhead view", where selected class is the category name. We also use the same time difference embedding network and then fine tune the stable diffusion 1.2. During inference, for both datsets we use 50 NFEs, and then perform optimization every 5 steps. More implementation details can be found in the supplementary materials. For WorldStrat, we evaluate our algorithm on first 1000 images in the validation set. For fMoW, we select the first 100 images in the validation set from 6 classes: Airport, Amusement Park, Car Dealership, Crop Field, Educational Institution, Electric Substation for evaluation.
@@
We consider six state-of-the-art baselines, WorldStrat [Cornebise et al., 2022], MSRResNet [Wang et al., 2018b], DBPN [Haris et al., 2018], Pix2Pix [Isola et al., 2017], ControlNet [Zhang et al., 2023], and DiffusionSat [Khanna et al., 2024]. Both ControlNet and DiffusionSat are diffusion-based methods that takes LR as an additional condition for fine-tuned stable diffusion to predict the HR image. Both MSRResNet and DBPN are CNN-based methods that directly map the low-resolution image to the high-resolution image, and Pix2Pix is a GAN-based method. Both WorldStrat and DiffusionSat are fusion-based methods that fuse multiple low-resolution input to predict the HR image, while others take single low-resolution input and predict the HR image.
@@
ControlNet [Zhang et al., 2023] 1427 12500 1427 21400
DiffusionSat [Khanna et al., 2024] 1428 12500 1428 20000
@@ 5 Conclusion
Appendix A Appendix
@@ A.1 Implementation Details
WorldStrat We obtain the RGB satellite images provided by [Cornebise et al., 2022]. We follow the preprocessing steps given by [Cornebise et al., 2022], that crop the HR images into patches of $192\times 192$ and LR images into corresponding pathes of the HR images of size $63\times 63$. For each HR image, there are 16 corresponding LR images of different time. We extract the RGB band of the HR images and the RGB band of every LR images. We then resize LR image into the size of $192\times 192$. We use the same training and validation split provided by [Cornebise et al., 2022] and then form a training and validation set. We then extract the timestamp from the metadata and compute the $dt_{i}$ for each $\text{LR}_{i}$.
fMoW We obtain the high-resolution images from [Christie et al., 2018], and the paired Sentinel-2 images from [Cong et al., 2022]. We first identify the area of interest on the HR images as given by [Christie et al., 2018], and then crop out other areas. Then we crop the corresponding LR images following the pre-processing steps given by [Cong et al., 2022]. We crop the HR images into patches of $512\times 512$, and align LR images into patches based on each HR image patch. Then we resize each LR image into the size of $512\times 512$ in accordance with the HR image. We consider 6 categories: airport, amusement parks, car dealership, crop field, educational institution, electric substation for training and testing. For training, we filter out HR images that do not have a corresponding LR, and those do not have three channels. We take the same training and validation split from [Christie et al., 2018]. When training, we consider all images from [Christie et al., 2018], and when testing, we pick the first 100 images from each selected category of the validation set of [Christie et al., 2018]. We also extract $dt_{i}$ from the metadata of fmow provided by [Christie et al., 2018], and the metadata of Sentinel-2 data provided by [Cong et al., 2022].
@@
We take the pretrained checkpoint (SD1.2) provided by [Rombach et al., 2022], and then fine tune on the processed [Christie et al., 2018] and [Cornebise et al., 2022] datasets. We rescale every LR image and HR image to the scale of [0,1]. Then, we use a learning rate $1e^{-5}$, and a batch size of 4 for both datasets. We stop training when the FID of sampled images stops improving. For [Cornebise et al., 2022] dataset, we only train 800 iterations (partly due to the small dataset size). For [Christie et al., 2018] dataset, we train 4400 iterations. We then take the model for downstream inference tasks. For WorldStrat dataset, we use the prompt "Satellite images" for training. For fMoW dataset, we use the prompt "Satellite image of selected class from a direct overhead view", where selected class is the category name.
@@ A.3 Implementation Details of Baselines
We follow the original codebase of [Cornebise et al., 2022], where we train the HighResNet model on both [Christie et al., 2018] and [Cornebise et al., 2022] datasets. We tune the hyperparameters of the loss function based on our validation set. We stop training when the validation performance converges. On [Cornebise et al., 2022] we train 125000 iterations with a batch size of 32, and on [Christie et al., 2018] we take 100000 iterations with a batch size of 32.
@@
We follow the original codebase of [Wang et al., 2018b], where we train the MSRResNet model on both [Christie et al., 2018] and [Cornebise et al., 2022] datasets. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 160000 iterations for both [Cornebise et al., 2022] and [Christie et al., 2018] datasets with a batch size of 16.
@@
We follow the original codebase of [Haris et al., 2018], where we train the MSRResNet model on both [Christie et al., 2018] and [Cornebise et al., 2022] datasets. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 100 epochs for both [Cornebise et al., 2022] and [Christie et al., 2018] datasets with a batch size of 16.
@@
We follow the original codebase of [Isola et al., 2017], where we train the MSRResNet model on both [Christie et al., 2018] and [Cornebise et al., 2022] datasets. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 30 epochs for [Christie et al., 2018] and 100 epochs for [Cornebise et al., 2022] with a batch size of 16.
@@
We follow the original codebase of [Zhang et al., 2023]. During training, we randomly pick a LR image and its paired HR image. We tune the hyperparameters of the loss function based on validation set performance. We train for 12500 iterations with a batch size of 16 for [Cornebise et al., 2022], and 21500 iterations with a batch size of 16 for [Christie et al., 2018].
@@
We implemented the 3D ControlNet architecture as mentioned in [1]. Then, we take the RGB band and the SWIR, NIR band from LR image for training 3D ControlNet. We tune the hyperparameters of the loss function based on validation set performance. We train for 12500 iterations with a batch size of 16 for [Cornebise et al., 2022], and 20000 iterations with a batch size of 16 for [Christie et al., 2018]. We observe that further training worsened FID scores on both datasets.
@@ A.4 More Results
WorldStrat [Cornebise et al., 2022] 17.98 0.396 13.42 0.443
MSRResNet [Wang et al., 2018b] 19.81 0.512 13.01 0.290
DBPN [Haris et al., 2018] 19.17 0.471 11.90 0.268
Pix2Pix [Isola et al., 2017] 19.76 0.448 12.21 0.180
ControlNet [Zhang et al., 2023] 11.89 0.113 10.82 0.117
DiffusionSat [Khanna et al., 2024] 12.34 0.133 10.63 0.109